The length of a pencil is measured to be 10.00 cm with a ruler. What is the fractional error in the length of the pencil?

0.05%

0.005

0.005 cm

0.005 mm

None of the above.
Answer #1

Least count with which measurement has been done = 0.01 cm

That means ∆L = 0.01 cm

So fractional error = ∆L/L = 0.01/10.00 = 0.0001

So none of the above